OpenStudy (anonymous):

how to change 490x2 + 90y2 - 4410=0 into standard form?

OpenStudy (nikita2):

490x2 + 90y2 - 4410=0 490x2 + 90y2 = 4410 490/4410x2 + 90/4410y2 = 1 49/441x2 + 9/441y2 = 1

OpenStudy (nikita2):

x^2/9 + y^2/49 = 1